RITZ ASSIGNMENT AND AWAT ANALYSIS OF THE RING-PUCKERING HOT BANDS OF OXETANE

\underline{GIOVANNI MORUZZI}, Dipartimento di Fisica dell’Università di Pisa and INFM, Via Filippo Buonarroti 2, I-56127 Pisa, Italy; MARC KUNZMANN, Institut für Physikalische Chemie der Georg-August-Universität Göttingen, Germany; BRENDA P. WINNEWISSER AND MANFRED WIN-NEWISSER, Physikalisch-Chemisches Institut, Justus-Liebig-Universität, Heinrich-Buff-Ring 58, D-35392 Gießen, Germany.

The oxetane molecule consists of 10 atoms, corresponding to 24 vibrational degrees of freedom. These include one largeamplitude vibration, i.e., the ring puckering, with its fundamental at approximately 53 cm^{−1}. The first rotationally resolved spectrum of these transitions has been reported recently. We have investigated the oxetane spectrum between 52 and 162 cm^{−1}. Six different Δv_{rp} = 1 rotation-puckering transitions are observed in this region. Preliminary analyses of the lowest three bands have already been presented^{a}^{b}^{c}. In the present contribution we complete our assignment work in the investigated spectral region. We shall present the results of fitting our FIR assignments and previously published MW and MMW data with the AWAT1 program (Watson Hamiltonian for an asymmetric rotor, A-reduction, programmed by K. M. T. Yamada), and with an equivalent program now incorporated into the Ritz program, for levels up to v_{rp} = 5. Watsonian frequency predictions have been used as a feed-back for the assignments in the most dense and entangled spectral regions. An analysis of the dependence of the rotational constants and of the quartic centrifugal distortion constants of the Watsonian on the puckering state will also be attempted.

aM. Winnewisser, M. Kunzmann and M. Lock, 53rd Ohio State University International Symposium on Molecular Spectroscopy, Columbus, June 15–19, 1998, TD04

bG. Moruzzi, M. Kunzmann, B. P. Winnewisser and M. Winnewisser, 15th International Conference on High Resolution Molecular Spectroscopy, Prague, Czech Republic, August 30 – September 3, 1998, H7

cG. Moruzzi, M. Kunzmann, B. P. Winnewisser and M. Winnewisser, 16th International Colloquium on High Resolution Molecular Spectroscopy, Dijon, France, September 6–10, 1999, F20